I JUST NEED ANSWER FOR "C" TO MAKE SURE I'M CORRECT PLEASE

Yuki (age 45 at year-end) has been contributing to a traditional IRA for years (all deductible contributions) and her IRA is now worth $46,000. She is trying to decide whether she should roll over her traditional IRA into a Roth IRA. Her current marginal tax rate is 25 percent. She plans to withdraw the entire balance of the account in 20 years and she expects to earn a before-tax rate of return of 5.6 percent on her retirement accounts and a 4.6 percent after-tax rate of return on all investments outside of her retirement accounts For each of the following alternative scenarios, indicate how much more or less Yuki will accumulate after taxes in 20 years if she rolls over her traditional IRA into a Roth IRA. Be sure to include the opportunity cost of having to pay taxes on the rollover. (Round Future value factor to 4 decimal places. Round final answers to the nearest whole dollar amount.) a.When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 35 percent. Accumulation if she keeps funds in traditional IRA Accumulation if she rolls funds into Roth IRA Greater $ 88,910 S 108,514 19,604 accumulation if she rolls over traditional IRA into Roth IRA b.When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 20 percent. Accumulation if she keeps funds in traditional IRA Accumulation if she rolls funds into Roth IRA $ 109,427 S 108,514 S sser accumulation if she rolls over traditional IRA into Roth IRA c.Assume the same facts as in part (b), except that she earns a 3.6 percent after-tax rate of return on investments outside of the retirement accounts Accumulation if she keeps funds in traditional IRA Accumulation if she rolls funds into Roth IRA Greater accumulation if she rolls over traditional IRA into Roth IRA
Answer #1

Part C

Accumulation if she keeps funds in traditional IRA = 46000*(1.056^20) *(1-20%) = $109427

Accumulation if she rolls funds into Roth IRA = (46000*(1.056^20))-(46000*25%*(1.036^20)) = $113455

Difference = 113455-109427 = $4028

Therefore,

Accumulation if she keeps funds in traditional IRA

$109427

Accumulation if she rolls funds into Roth IRA

$113455

Greater

Accumulation if she rolls over traditional IRA into Roth IRA

$4028